# High-quality Surface Reconstruction using Gaussian Surfels

### v1

@@ Abstract.
We propose a novel point-based representation, Gaussian surfels, to combine the advantages of the flexible optimization procedure in 3D Gaussian points and the surface alignment property of surfels. This is achieved by directly setting the z-scale of 3D Gaussian points to 0, effectively flattening the original 3D ellipsoid into a 2D ellipse. Such a design provides clear guidance to the optimizer. By treating the local z-axis as the normal direction, it greatly improves optimization stability and surface alignment. While the derivatives to the local z-axis computed from the covariance matrix are zero in this setting, we design a self-supervised normal-depth consistency loss to remedy this issue. Monocular normal priors and foreground masks are incorporated to enhance the quality of the reconstruction, mitigating issues related to highlights and background. We propose a volumetric cutting method to aggregate the information of Gaussian surfels so as to remove erroneous points in depth maps generated by alpha blending. Finally, we apply screened Poisson reconstruction method to the fused depth maps to extract the surface mesh. Experimental results show that our method demonstrates superior performance in surface reconstruction compared to state-of-the-art neural volume rendering and point-based rendering methods.
@@ 3.5. Implementation details
We support the use of sparse points computed via the Structure from Motion (SfM) as the initialization. We found this could accelerate the decrease of loss function values in the first few steps but overall did not significantly improve the convergence rate. Thus, we initialize our Gaussian surfels with random positions and rotations inside the approximated bounding box of the target object in all our experiments.
@@ 4. Experiments
Table 1. Geometry quality comparison. We report the Chamfer distance (mm) and the averaged training time (minute) on DTU dataset compared with baselines. Best results are highlighted as 1st, 2nd, and 3rd.
Scans Ours 3DGS SuGaR NeuS NeuS2 INSR
24 0.90 2.89 1.85 0.83 0.56 2.86
37 1.25 2.65 1.19 0.98 0.76 2.81
40 1.12 2.66 1.91 0.56 0.49 2.09
55 0.68 2.31 1.64 0.37 0.37 0.81
63 1.16 3.55 2.76 1.13 0.92 1.65
65 1.53 2.94 1.94 0.59 0.71 1.39
69 1.07 2.19 1.97 0.60 0.76 1.47
83 1.88 2.66 3.07 1.45 1.22 1.67
97 1.97 2.99 2.38 0.95 1.08 2.47
105 0.98 1.83 1.33 0.78 0.63 1.12
106 1.17 2.54 1.69 0.52 0.59 1.22
110 1.78 3.38 3.61 1.43 0.89 2.30
114 0.60 2.05 1.53 0.36 0.40 0.98
118 0.93 1.99 1.83 0.45 0.48 1.41
122 0.85 2.02 1.97 0.45 0.55 0.95
Mean 1.19 2.58 2.05 0.77 0.70 1.68
Time 6.67 5.19 30.9 408 3.27 8.48
Table 2. Geometry quality comparison on BlendedMVS dataset. Best results are highlighted as 1st, 2nd, and 3rd.
Methods Ours 3DGS SuGaR NeuS NeuS2 INSR
Stone 4.45 9.59 11.1 5.04 4.55 4.87
Clock 3.19 2.95 4.82 3.61 3.28 4.47
Dog 1.94 5.53 8.32 1.94 2.05 2.71
Durian 7.13 11.4 11.5 7.48 8.39 10.7
Man 2.56 6.72 8.73 1.97 1.91 1.93
Bear 4.72 6.16 9.81 4.05 3.95 2.92
Jade 1.35 1.28 3.78 1.31 0.99 1.20
Mean 3.62 6.32 8.29 3.63 3.59 4.12
Time 3.82 3.47 18.3 377 3.61 2.78
Figure 7. Qualitative comparisons on DTU and BlendedMVS Datasets.
@@
We first benchmark our method on the DTU dataset (Jensen et al., 2014) both quantitatively (Table 1) and qualitatively (Fig. 7). The dataset encompasses laboratory-captured scenes, each encompassing 49 or 64 images with resolution $1600\times 1200$, and we choose the same set of scenes selected by IDR (Yariv et al., 2020) that contain 15 objects with manually annotated object masks. We further tested on 7 challenging scenes from the low-res set of the BlendedMVS dataset (Yao et al., 2020), where each scene comprises 24 to 143 images at $768\times 576$ pixels with masks provided. The evaluation result is obtained using the DTU evaluation code.
The statistics of Chamfer distances in these two tests are shown in Tables 1 and 2. Our method significantly outperforms 3DGS and SuGaR on the DTU and the BlendedMVS datasets. Compared with NeuS2, our results show a larger Chamfer distance. We hypothesize that it is due to the bias inherent in per-view depth calculation using alpha blending. Despite the higher Chamfer distance, our results often have less noise than NeuS2 and more details than NeuS, as illustrated in Fig. 7. In addition, our method can be applied to the reconstruction of open surfaces (Fig. 2), since Gaussian surfels do not assume closed surfaces as in the representation of signed distance functions. Overall, our method achieves a good balance between reconstruction speed and quality. In contrast to NeuS, our approach exhibits rapid convergence to high-quality reconstructions, as detailed in Table 1 and 2. Furthermore, while NeuS (Wang et al., 2021) employs a large MLP to model surfaces, offering robustness to outliers, it may result in over-smoothing of the reconstructed surfaces as depicted in Fig. 7. On the other hand, our reconstruction speed is comparable to that of INSR (Zhao et al., 2022) and NeuS2 (Wang et al., 2023), both of which utilize hash feature maps and tiny MLPs to expedite implicit surface optimization.
@@
We show quantitative analyses of our loss terms by excluding each individually from the optimization (Table 4). The photometric loss $\mathcal{L}_{\mathrm{p}}$ is always included. In addition, we evaluate the effect of volumetric cutting by excluding volumetric cutting (w/o cut). Removing depth-normal consistency loss $\mathcal{L}_{\mathrm{c}}$ significantly reduces reconstruction quality. The qualitative results of the ablation are shown in Fig. 8. The normal-prior loss helps reduce fluctuations in the geometric reconstruction in highlighted areas (the second row of Fig. 8), contributing to a lower Chamfer distance. However, as indicated in the third colomn of Table 4, removing the normal-prior loss can slightly improve the rendering quality. This is because this loss might prevent the optimizer from approximating the highlights as a virtual light sources behind the actual surface, resulting in lower rendering quality. The opacity loss aids in representing details (the fourth column in 4). As depicted in the third rows of Fig. 8, with this loss, the details on the “LEGO studs” are much more clearly reconstructed and rendered.
The last two columns of Table 4 demonstrate the efficacy of mask loss $\mathcal{L}_{\mathrm{m}}$ and volumetric cutting in noise reduction, which can improve the quality of reconstruction. Fig. 10 presents qualitative results without masks, where the absence of masks may lead to noises in texture-less background regions.
Figure 8. Top: Ablation on depth-normal consistency loss. For the normal column, the part left to the central dashed line on the object represents the rendered normals, while the right part represents the normals calculated from the rendered depth. We can observe better consistency with the depth-normal consistency loss, resulting in improved reconstruction. Middle: Ablation on normal-prior loss; the concave parts on the surface caused by highlights are corrected. Bottom: Ablation on opacity loss. The LEGO studs are better reconstructed, with sharper edges.
@@ 5. Conclusion
Figure 10. Visualization of our method trained with and without foreground mask on DTU-#83.

### v2

@@ Abstract.
We propose a novel point-based representation, Gaussian surfels, to combine the advantages of the flexible optimization procedure in 3D Gaussian points and the surface alignment property of surfels. This is achieved by directly setting the z-scale of 3D Gaussian points to 0, effectively flattening the original 3D ellipsoid into a 2D ellipse. Such a design provides clear guidance to the optimizer. By treating the local z-axis as the normal direction, it greatly improves optimization stability and surface alignment. While the derivatives to the local z-axis computed from the covariance matrix are zero in this setting, we design a self-supervised normal-depth consistency loss to remedy this issue. Monocular normal priors and foreground masks are incorporated to enhance the reconstruction quality, mitigating issues related to highlights and background. We propose a volumetric cutting method to aggregate the information of Gaussian surfels so as to remove erroneous points in depth maps generated by alpha blending. Finally, we apply screened Poisson reconstruction method to the fused depth maps to extract the surface mesh. Experimental results show that our method demonstrates superior performance in surface reconstruction compared to state-of-the-art neural volume rendering and point-based rendering methods.
@@ 3.5. Implementation details
We support the use of sparse points computed via the Structure from Motion (SfM) as the initialization (Fig.  10). We found this could accelerate the decrease of loss function values in the first few steps but overall did not significantly improve the convergence rate. Thus, we initialize our Gaussian surfels with random positions and rotations inside the approximated bounding box of the target object in all our experiments.
@@ 4. Experiments
Table 1. Geometry quality comparison on the DTU dataset. We report the Chamfer distance compared with baselines. Mean scores of Chamfer distance (mm) and training times (minute) are included on the bottom. Best results are highlighted as 1st, 2nd, and 3rd.
Methods Ours 3DGS SuGaR NeuS NeuS2 INSR
24 0.66 2.89 1.85 0.83 0.56 2.86
37 0.93 2.65 1.19 0.98 0.76 2.81
40 0.54 2.66 1.91 0.56 0.49 2.09
55 0.41 2.31 1.64 0.37 0.37 0.81
63 1.06 3.55 2.76 1.13 0.92 1.65
65 1.14 2.94 1.94 0.59 0.71 1.39
69 0.85 2.19 1.97 0.60 0.76 1.47
83 1.29 2.66 3.07 1.45 1.22 1.67
97 1.53 2.99 2.38 0.95 1.08 2.47
105 0.79 1.83 1.33 0.78 0.63 1.12
106 0.82 2.54 1.69 0.52 0.59 1.22
110 1.58 3.38 3.61 1.43 0.89 2.30
114 0.45 2.05 1.53 0.36 0.40 0.98
118 0.66 1.99 1.83 0.45 0.48 1.41
122 0.53 2.02 1.97 0.45 0.55 0.95
Mean 0.88 2.58 2.05 0.77 0.70 1.68
Time 6.67 5.19 30.9 408 3.27 8.48
Table 2. Geometry quality comparison on the BlendedMVS dataset. Best results are highlighted as 1st, 2nd, and 3rd.
Methods Ours 3DGS SuGaR NeuS NeuS2 INSR
Basketball 1.55 5.59 8.00 2.96 2.48 2.67
Bear 1.96 5.08 9.73 3.00 3.20 3.72
Bread 1.17 3.98 7.65 2.85 2.47 2.76
Camera 2.34 5.85 7.77 2.61 2.89 3.25
Clock 3.41 7.34 9.21 2.75 2.87 3.45
Cow 2.17 5.38 8.69 2.04 2.29 2.11
Dog 2.89 7.16 9.27 2.75 2.97 2.54
Doll 2.13 6.25 8.75 2.17 2.23 2.44
Dragon 2.75 4.90 9.76 2.95 2.68 2.20
Durian 2.35 6.66 8.04 3.14 3.39 5.78
Fountain 3.01 5.03 9.05 3.03 2.68 3.79
Gundam 0.85 4.46 7.32 1.62 1.87 1.36
House 2.12 4.64 7.28 3.23 3.02 2.93
Jade 3.50 6.96 10.7 4.25 4.03 3.85
Man 2.47 7.13 9.29 2.29 2.20 2.22
Monster 1.38 5.84 8.20 1.92 1.96 1.80
Sculpture 2.90 6.68 8.98 2.10 2.06 1.70
Stone 1.84 6.01 9.19 2.51 2.01 2.12
Mean 2.27 5.83 8.71 2.67 2.63 2.82
Time 3.82 3.47 18.3 377 3.61 2.78
@@
We first benchmark our method on the DTU dataset (Jensen et al., 2014) both quantitatively (Table 1) and qualitatively (Fig. 8). The dataset encompasses laboratory-captured scenes, each encompassing 49 or 64 images with resolution $1600\times 1200$, and we choose the same set of scenes selected by IDR (Yariv et al., 2020) that contain 15 objects with manually annotated object masks. We further tested on 18 challenging scenes from the low-res set of the BlendedMVS dataset (Yao et al., 2020), where each scene comprises 24 to 143 images at $768\times 576$ pixels with masks provided. The evaluation result is obtained using the DTU evaluation code.
The statistics of Chamfer distances in these two tests are shown in Tables 1 and 2. Our method significantly outperforms 3DGS and SuGaR on the DTU and the BlendedMVS datasets. Compared with NeuS2, our results show a larger Chamfer distance on DTU. We hypothesize that it is due to the bias inherent in per-view depth calculation using alpha blending. Despite the higher Chamfer distance, our results often have less noise than NeuS2 and more details than NeuS, as illustrated in Fig. 8. In addition, our method can be applied to the reconstruction of open surfaces (Fig. 2), since Gaussian surfels do not assume closed surfaces as in the representation of signed distance functions. Overall, our method achieves a good balance between reconstruction speed and quality. In contrast to NeuS, our approach exhibits rapid convergence to high-quality reconstructions, as detailed in Table 1 and 2. Furthermore, while NeuS (Wang et al., 2021) employs a large MLP to model surfaces, offering robustness to outliers, it may result in over-smoothing of the reconstructed surfaces as depicted in Fig. 8. On the other hand, our reconstruction speed is comparable to that of INSR (Zhao et al., 2022) and NeuS2 (Wang et al., 2023), both of which utilize hash feature maps and tiny MLPs to expedite implicit surface optimization.
@@
Figure 9. Top: Ablation on depth-normal consistency loss. For the normal column, the part left to the central dashed line on the object represents the rendered normals, while the right part represents the normals calculated from the rendered depth. We can observe better consistency with the depth-normal consistency loss, resulting in improved reconstruction. Second row: Ablation on normal-prior loss; the concave parts on the surface caused by highlights are corrected. Third row: Ablation on opacity loss. The LEGO studs are better reconstructed, with sharper edges. Bottom Ablation without masks, where the absence of masks may lead to noises in texture-less background regions.
@@
We show quantitative analyses of our loss terms by excluding each individually from the optimization (Table 4). The photometric loss $\mathcal{L}_{\mathrm{p}}$ is always included. In addition, we evaluate the effect of volumetric cutting by excluding volumetric cutting (w/o cut). Removing depth-normal consistency loss $\mathcal{L}_{\mathrm{c}}$ significantly reduces reconstruction quality. The qualitative results of the ablation are shown in Fig. 9. The normal-prior loss helps reduce fluctuations in the geometric reconstruction in highlighted areas (the second row of Fig. 9), contributing to a lower Chamfer distance. However, as indicated in the third colomn of Table 4, removing the normal-prior loss can slightly improve the rendering quality. This is because this loss might prevent the optimizer from approximating the highlights as a virtual light sources behind the actual surface, resulting in lower rendering quality. The opacity loss aids in representing details (the fourth column in 4). As depicted in the third rows of Fig. 9, with this loss, the details on the “LEGO studs” are much more clearly reconstructed and rendered. The last two columns of Table 4 demonstrate the efficacy of mask loss $\mathcal{L}_{\mathrm{m}}$ and volumetric cutting in noise reduction, which can improve the quality of reconstruction.
Figure 10. Reconstruction results of our method on indoor scenes from DeepBlending (Hedman et al., 2018) and MipNeRF360 (Barron et al., 2022) datasets when initialized with sparse SfM points.